# Matching Markets Meet LLMs: Algorithmic Reasoning with Ranked Preferences

### v1

@@ I.2.2 CoT-Shortlist
The main distinction between the prompt described here and the one in Section I.2.1 lies in how the algorithm’s execution is detailed. In the CoT-Shortlist prompt, the provided algorithm trace includes an additional step: agents remove each other from their respective shortlists if they become matched with a partner they find more desirable than the other agents on their list. All other aspects of the prompt are identical to the CoT-Vanilla prompt.
[…]
I.2.3 Few-shot Examples
@@
Few-shot prompting was applied to the prompt in Section I.1 by additionally including a series of randomly generated preference/stable solution pairs. As with other few-shot prompting strategies, the model is then asked to generate a stable solution (as shown in Section I.1). As with the CoT methods, each sample preference/stable solution pairs is enclosed in $<$example$>$ XML tags.
[…]
I.3 Example Prompt for Evaluating Stability
The following prompt requires LLMs to determine if a given solution to a provided preference profile is stable. Unlike the prompt in Section I.1, the only element that the LLM must include in their response is a binary response (yes/no).
@@ I.3 Example Prompt for Evaluating Stability
Consider the following instance of the two-sided matching problem, where 5 men are to be matched with 5 women. Here are the preference lists for all individuals: <<<preferences>>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], }} <<</preferences>>> Your task is to dermine whether the following matching is stable or not. <<<matching>>> [[M1, W4],[M2, W5],[M3, W3],[M4, W1],[M5, W2],] <<</matching>>> Please return ’Yes’ if you think the provided matching is stable and ’No’ if you think it is unstable, and enclose your answer in <<<answer>>><<</answer>>> tags.
I.4 Example Prompts for Preference Comprehension
@@
In each of the following preference comprehension prompts, models are asked to provide the name of an agent (in level-1) or to provide a binary answer (yes/no for levels 2 and 3) in response to a provided question. In addition to changing the preference profiles for each instance of a preference comprehension task, the agents and positions mentioned in the question are also changed with each instance. For details about each level of preference comprehension, view Section 6.
I.4.1 Level-1
@@
Your goal is to correctly interpret the given preference lists and answer a specific question about agent preferences. First, here are the preference lists for all individuals: <<<preferences>>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], W2: [M1,M3,M4,M5,M2], }} <<</preferences>>> Now, you will be asked a specific question about agent preferences: <<<question>>> Who is agent W2’s, 1-most preferred agent? <<</question>>> Once you have determined the answer, provide your output in the following format: 1. The solution as a single agent name. For example, "W1" Present your final answer within <<<answer>>> tags. IMPORTANT: ONLY RETURN THE NAME OF THE SINGLE AGENT THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
I.4.2 Level-2
@@
You are an AI assistant tasked with analyzing preference profiles in a two-sided matching problem with one-to-one solutions. Your goal is to correctly interpret the given preference lists and answer a specific question about agent preferences. First, here are the preference lists for all individuals: <<<preferences>>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], }} <<</preferences>>> Now, you will be asked a specific question about agent preferences: <<<question>>> Would agent W1, prefer M3 or M2 over M4? <<</question>>> Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within <<<answer>>> tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
I.4.3 Level-3
@@
You are an AI assistant tasked with analyzing preference profiles in a two-sided matching problem with one-to-one solutions. Your goal is to correctly interpret the given preference lists and answer a specific question about agent preferences. First, here are the preference lists for all individuals: <<<preferences>>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], }} <<</preferences>>> Now, you will be asked a specific question about agent preferences: <<<question>>> If agent W1 is currently engaged to M4, would she accept proposals from M3 or M2? <<</question>>> Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within <<<answer>>> tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
I.5 Example Prompt for Resolving Instability
@@
For the task of resolving instability in a given unstable solution, the prompt begins by providing models with the instance’s preference profile (as with the prompts for the other tasks). In addition, the prompt includes an unstable matching, and asks the model to resolve the instability by outputting a stable solution (in an identical format to the prompt in Section I.1).
@@
You are an intelligent assistant who is an expert in algorithms. Consider the following instance of the two-sided matching problem and respective unstable matching, where 5 men are to be matched with 5 women. Here are the preference lists for all individuals: <<<preferences>>> { M: { M1: [W5,W3,W4,W2,W1], M2: [W2,W3,W5,W1,W4], M3: [W5,W3,W1,W4,W2], M4: [W1,W3,W2,W5,W4], M5: [W2,W3,W4,W1,W5], }, W: { W1: [M3,M5,M4,M1,M2], W2: [M1,M3,M4,M5,M2], W3: [M3,M2,M4,M1,M5], W4: [M4,M2,M3,M5,M1], W5: [M2,M4,M5,M1,M3], }} <<</preferences>>> Here is an unstable matching. <<<answer>>> { "M1": "W4", "M2": "W5", "M3": "W3", "M4": "W2", "M5": "W1" } <<</answer>>> Your task is to modify the given unstable matching to make it equivalent to the proposer-optimal stable matching. You can use XML tags like <<<scratchpad>>> to explain your thought process while computing the solution. Once you have found a stable matching, please return your matching in the JSON format given below: <<<answer>>> { "M1": "<<<woman matched with M1>>>", "M2": "<<<woman matched with M2>>>", "M3": "<<<woman matched with M3>>>", "M4": "<<<woman matched with M4>>>", "M5": "<<<woman matched with M5>>>" } <<</answer>>> Make sure that each man/woman is matched with exactly ONE partner. It is mandatory that you provide a matching as a JSON object enclosed in <<<answer>>><<</answer>>> tags as described above.
I.6 Example Prompt for Repeated Queries Due to Missing JSON Object
@@
For tasks where the desired output is a JSON object (when outputting a stable solution), models are given an additional opportunity to rectify issues in their response if the original response is incorrectly formatted. The prompt below is passed to the model to help rectify issues related to missing JSON objects. Note that the $<$initially passed prompt$>$ and $<$last 3,000 characters of LLM’s first response$>$ XML tags are replaced by the initial prompt and the tail of the models initial response, respectively.
@@
Previously, I gave you the following task: ——————————————————— <<<initially passed prompt>>> ——————————————————— In your response, you either failed to provide me with a matching or did not adhere to the JSON format I had asked for. Here are the last few lines of your response for reference: ——————————————————— <<<last 3,000 characters of LLM’s first response>>> ——————————————————— Please correct your response and provide me with the matching in the following JSON format, enclosed in <<<answer>>><<</answer>>> tags.<<<answer>>> { "M1": "<<<woman matched with M1>>>", "M2": "<<<woman matched with M2>>>", "M3": "<<<woman matched with M3>>>", "M4": "<<<woman matched with M4>>>", "M5": "<<<woman matched with M5>>>" } <<</answer>>> Make sure that each man/woman is matched with exactly ONE partner.
@@
I.7 Example Prompt for Repeated Queries Due to Incorrectly Formatted JSON Object
@@
Similar to the prompt in Section I.6, the following prompt is passed to the model when the model’s initial response contains an incorrectly formatted JSON object. Once again, the $<$initially passed prompt$>$ and $<$last 3,000 characters of LLM’s first response$>$ XML tags are replaced by the initial prompt and the tail of the models initial response, respectively.
@@
Previously, I gave you the following task: ——————————————————— <<<initially passed prompt>>> ——————————————————— In your response, you failed adhere to the JSON format I had asked for. Here are the last few lines of your response for reference: ——————————————————— <<<last 3,000 characters of LLM’s first response>>> ——————————————————— Please correct your response and provide me with the matching in the following JSON format, enclosed in <<<answer>>><<</answer>>> tags.<<<answer>>> { "M1": "<<<woman matched with M1>>>", "M2": "<<<woman matched with M2>>>", "M3": "<<<woman matched with M3>>>", "M4": "<<<woman matched with M4>>>", "M5": "<<<woman matched with M5>>>" } <<</answer>>> Make sure that each man/woman is matched with exactly ONE partner.
I.8 Example Prompt for Repeated Queries Due to Incomplete Matching
@@
Similar to the prompt in Section I.6, the following prompt is passed to the model when the model’s initial response contains a correctly formatted JSON object, but the matching itself is incomplete, or some agents have multiple partners. After the initially passed prompt, note that additional details are provided to assist the LLM in rectifying its response. Again, the $<$initially passed prompt$>$ and $<$last 3,000 characters of LLM’s first response$>$ XML tags are replaced by the initial prompt and the tail of the models initial response, respectively.
@@
Previously, I gave you the following task: ——————————————————— <<<initially passed prompt>>> ——————————————————— In your response, the matching you selected involves some women being matched with multiple men, which is not allowed. For example, W2 is matched with M1, M2, and M5. Additionally, W3, and W4 are unmatched. Here are the last few lines of your response for reference: ——————————————————— <<<last 3,000 characters of LLM’s first response>>> ——————————————————— Please correct your response and provide me with the matching in the following JSON format, enclosed in <<<answer>>><<</answer>>> tags.<<<answer>>> { "M1": "<<<woman matched with M1>>>", "M2": "<<<woman matched with M2>>>", "M3": "<<<woman matched with M3>>>", "M4": "<<<woman matched with M4>>>", "M5": "<<<woman matched with M5>>>" } <<</answer>>> Make sure that each man/woman is matched with exactly ONE partner.
Appendix J Training Examples for Fine-tuning
J.1 System-prompt ($\boldsymbol{s}$)
@@
J.2 High-level instruction ($\boldsymbol{u}$)
@@
J.3 Preference Profile ($\boldsymbol{p^{(i)}}$)
<<<preferences>>> { M: { M1: [W5,W3,W4,W2,W1], M2: [W2,W3,W5,W1,W4], M3: [W5,W3,W1,W4,W2], M4: [W1,W3,W2,W5,W4], M5: [W2,W3,W4,W1,W5], }, W: { W1: [M3,M5,M4,M1,M2], W2: [M1,M3,M4,M5,M2], W3: [M3,M2,M4,M1,M5], W4: [M4,M2,M3,M5,M1], W5: [M2,M4,M5,M1,M3], }} <<</preferences>>> <<</preferences>>>
J.4 Task-prompt ($\boldsymbol{t^{(i)}}$)
@@
    Your task is to find the proposer-optimal stable matching. Once you have found a stable matching, please return your matching in the JSON format given below: <<<answer>>> { "M1": "<<<woman matched with M1>>>", "M2": "<<<woman matched with M2>>>", "M3": "<<<woman matched with M3>>>", "M4": "<<<woman matched with M4>>>", "M5": "<<<woman matched with M5>>>" } <<</answer>>> Make sure that each man/woman is matched with exactly ONE partner. It is important that you enclose your JSON object in <<<answer>>><<</answer>>> tags.
@@
    Now, you will be asked a specific question about agent preferences: $<$question$>$ Who is agent W3’s, 5-most preferred agent? $<$/question$>$ Once you have determined the answer, provide your output in the following format: 1. The solution as a single agent name. For example, "W1" Present your final answer within $<$answer$>$ tags. IMPORTANT: ONLY RETURN THE NAME OF THE SINGLE AGENT THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
@@
    Now, you will be asked a specific question about agent preferences: $<$question$>$ Would agent W3, prefer M4 and M3 over M1? $<$/question$>$ Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within $<$answer$>$ tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
@@
    Now, you will be asked a specific question about agent preferences: $<$question$>$ If agent W5 is currently engaged to M3, would she accept proposals from M1 and M4? $<$/question$>$ Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within $<$answer$>$ tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
@@
### Response: $<$think$>$
J.5 Reasoning-trace ($\boldsymbol{r^{(i)}}$):
@@
    ### Response: $<$think$>$ Okay, so I need to find a stable matching for 5 men and 5 women. For this, I can use the Gale-Shapley algorithm. Before I compute the solution for this instance, I will list the steps of the Gale-Shapley algorithm. 1. Initialize all men and women as unmatched. 2. Create a list to keep track of each man’s next proposal (initially set to 0 for all men). 3. While there are unmatched men: a. Select an unmatched man (M). b. Find the next woman (W) on M’s preference list that he hasn’t proposed to yet. c. If W is unmatched, match M and W. d. If W is matched but prefers M to her current partner: - Unmatch W from her current partner. - Match M and W. - Set the unmatched man as W’s previous partner. e. If W rejects M, move to the next woman on M’s preference list. 4. Repeat step 3 until all men are matched. Next, I will exectute the above algorithm for the given instance. The steps are as follows: M1 is free. M1 proposes to W3 Since W3 is free, W3 accepts the proposal. Now M1 and W3 are matched. M2 is free. M2 proposes to W5 Since W5 is free, W5 accepts the proposal. Now M2 and W5 are matched. M3 is free. M3 proposes to W1 Since W1 is free, W1 accepts the proposal. Now M3 and W1 are matched. M4 is free. M4 proposes to W4 Since W4 is free, W4 accepts the proposal. Now M4 and W4 are matched. M5 is free. M5 proposes to W4 Since W4 prefers their current partner M4 to M5, W4 rejects the proposal. M4 and W4 are still matched, and M5 is still free. M5 is free. M5 proposes to W1 Since W1 prefers M5 to their current partner M3, W1 accepts the proposal. Now M5 and W1 are matched, and M3 is free. M3 is free. M3 proposes to W4 Since W4 prefers their current partner M4 to M3, W4 rejects the proposal. M4 and W4 are still matched, and M3 is still free. M3 is free. M3 proposes to W3 Since W3 prefers their current partner M1 to M3, W3 rejects the proposal. M1 and W3 are still matched, and M3 is still free. M3 is free. M3 proposes to W2 Since W2 is free, W2 accepts the proposal. Now M3 and W2 are matched. $<$/think$>$
@@
    Alright, I have been given the task of determining W3’s 5-most preferred agent. For this, let me look at her preference list in order: position 1: M4 position 2: M5 position 3: M1 position 4: M3 position 5: M2 Hence, W3’s 5-most preferred agent is M2. $<$/think$>$

### v2

@@ Appendix I Inference Details
The main distinction between the prompt described here and the one in Section˜J.2.1 lies in how the algorithm’s execution is detailed. In the CoT-Shortlist prompt, the provided algorithm trace includes an additional step: agents remove each other from their respective shortlists if they become matched with a partner they find more desirable than the other agents on their list. All other aspects of the prompt are identical to the CoT-Vanilla prompt.
You are an intelligent assistant who is an expert in algorithms. Your task is to find the proposer-optimal stable matching, for the two-sided matching problem. Here is an example to demonstrate how you should proceed: <<example>> <<preferences>> { M: { M1: [W4,W3,W5,W2,W1], M2: [W5,W4,W3,W1,W2], M3: [W5,W4,W1,W2,W3], M4: [W5,W4,W2,W1,W3], M5: [W2,W4,W5,W3,W1], }, W: { W1: [M5,M2,M3,M4,M1], W2: [M3,M4,M5,M1,M2], W3: [M4,M1,M2,M5,M3], W4: [M5,M1,M4,M3,M2], W5: [M1,M4,M5,M3,M2], }} <</preferences>> M1 is free. M1 proposes to W4. W4 accepts the proposal. Now M1 and W4 are matched. W1 deletes M4, M3, M2 from her list. M4, M3, M2 delete W4 from their list. M2 is free. M2 proposes to W5. W5 accepts the proposal. Now M2 and W5 are matched. M3 is free. M3 proposes to W5. W5 accepts the proposal. Now M3 and W5 are matched. W5 prefers M3, so W5 breaks her engagement with M2. W3 deletes M2 from her list. M2 delete W5 from their list. M4 is free. M4 proposes to W5. W5 accepts the proposal. Now M4 and W5 are matched. W5 prefers M4, so W5 breaks her engagement with M3. W4 deletes M5, M3 from her list. M5, M3 delete W5 from their list. M5 is free. M5 proposes to W2. W2 accepts the proposal. Now M5 and W2 are matched. W5 deletes M1, M2 from her list. M1, M2 delete W2 from their list. <<answer>> { "M1": "W4", "M2": "W3", "M3": "W1", "M4": "W5", "M5": "W2" } <</answer>> <</example>> Consider the following instance of the two-sided matching problem, where 10 men are to be matched with 10 women $\ldots$
J.2.3 Few-shot Examples
@@
Few-shot prompting was applied to the prompt in Section˜J.1 by additionally including a series of randomly generated preference/stable solution pairs. As with other few-shot prompting strategies, the model is then asked to generate a stable solution (as shown in Section˜J.1). As with the CoT methods, each sample preference/stable solution pairs is enclosed in <<example>> XML tags.
@@
You are an intelligent assistant who is an expert in algorithms. Your task is to find the proposer-optimal stable matching, for the two-sided matching problem. Here is an example to demonstrate how you should proceed: <<example>> <<preferences>> { M: { M1: [W5,W3,W4,W2,W1], M2: [W3,W4,W1,W2,W5], M3: [W5,W1,W4,W2,W3], M4: [W3,W2,W5,W1,W4], M5: [W3,W4,W2,W1,W5], }, W: { W1: [M1,M4,M3,M5,M2], W2: [M2,M4,M5,M1,M3], W3: [M1,M2,M4,M5,M3], W4: [M3,M5,M1,M4,M2], W5: [M5,M3,M4,M2,M1], }} <</preferences>> <<answer>> { "M1": "W3", "M2": "W1", "M3": "W5", "M4": "W2", "M5": "W4" } <</answer>> <</example>> <<example>> <</example>> <<example>> <</example>> Consider the following instance of the two-sided matching problem, where 10 men are to be matched with 10 women $\ldots$
J.3 Example Prompt for Evaluating Stability
The following prompt requires LLMs to determine if a given solution to a provided preference profile is stable. Unlike the prompt in Section˜J.1, the only element that the LLM must include in their response is a binary response (yes/no).
Consider the following instance of the two-sided matching problem, where 5 men are to be matched with 5 women. Here are the preference lists for all individuals: <<preferences>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], }} <</preferences>> Your task is to determine whether the following matching is stable or not. <<matching>> [[M1, W4],[M2, W5],[M3, W3],[M4, W1],[M5, W2],] <</matching>> Please return ’Yes’ if you think the provided matching is stable and ’No’ if you think it is unstable, and enclose your answer in <<answer>><</answer>> tags.
J.4 Example Prompts for Preference Comprehension
@@
In each of the following preference comprehension prompts, models are asked to provide the name of an agent (in level-1) or to provide a binary answer (yes/no for levels 2 and 3) in response to a provided question. In addition to changing the preference profiles for each instance of a preference comprehension task, the agents and positions mentioned in the question are also changed with each instance. For details about each level of preference comprehension, view Section˜6.
J.4.1 Level-1
@@
Your goal is to correctly interpret the given preference lists and answer a specific question about agent preferences. First, here are the preference lists for all individuals: <<preferences>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], W2: [M1,M3,M4,M5,M2], }} <</preferences>> Now, you will be asked a specific question about agent preferences: <<question>> Who is agent W2’s, 1-most preferred agent? <</question>> Once you have determined the answer, provide your output in the following format: 1. The solution as a single agent name. For example, "W1" Present your final answer within <<answer>> tags. IMPORTANT: ONLY RETURN THE NAME OF THE SINGLE AGENT THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
J.4.2 Level-2
@@
You are an AI assistant tasked with analyzing preference profiles in a two-sided matching problem with one-to-one solutions. Your goal is to correctly interpret the given preference lists and answer a specific question about agent preferences. First, here are the preference lists for all individuals: <<preferences>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], }} <</preferences>> Now, you will be asked a specific question about agent preferences: <<question>> Would agent W1, prefer M3 or M2 over M4? <</question>> Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within <<answer>> tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
J.4.3 Level-3
@@
You are an AI assistant tasked with analyzing preference profiles in a two-sided matching problem with one-to-one solutions. Your goal is to correctly interpret the given preference lists and answer a specific question about agent preferences. First, here are the preference lists for all individuals: <<preferences>> { M: { M1: [W5,W3,W4,W2,W1], }, W: { W1: [M3,M5,M4,M1,M2], }} <</preferences>> Now, you will be asked a specific question about agent preferences: <<question>> If agent W1 is currently engaged to M4, would she accept proposals from M3 or M2? <</question>> Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within <<answer>> tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
J.5 Example Prompt for Resolving Instability
@@
For the task of resolving instability in a given unstable solution, the prompt begins by providing models with the instance’s preference profile (as with the prompts for the other tasks). In addition, the prompt includes an unstable matching, and asks the model to resolve the instability by outputting a stable solution (in an identical format to the prompt in Section˜J.1).
@@
You are an intelligent assistant who is an expert in algorithms. Consider the following instance of the two-sided matching problem and respective unstable matching, where 5 men are to be matched with 5 women. Here are the preference lists for all individuals: <<preferences>> { M: { M1: [W5,W3,W4,W2,W1], M2: [W2,W3,W5,W1,W4], M3: [W5,W3,W1,W4,W2], M4: [W1,W3,W2,W5,W4], M5: [W2,W3,W4,W1,W5], }, W: { W1: [M3,M5,M4,M1,M2], W2: [M1,M3,M4,M5,M2], W3: [M3,M2,M4,M1,M5], W4: [M4,M2,M3,M5,M1], W5: [M2,M4,M5,M1,M3], }} <</preferences>> Here is an unstable matching. <<answer>> { "M1": "W4", "M2": "W5", "M3": "W3", "M4": "W2", "M5": "W1" } <</answer>> Your task is to modify the given unstable matching to make it equivalent to the proposer-optimal stable matching. You can use XML tags like <<scratchpad>> to explain your thought process while computing the solution. Once you have found a stable matching, please return your matching in the JSON format given below: <<answer>> { "M1": "<<woman matched with M1>>", "M2": "<<woman matched with M2>>", "M3": "<<woman matched with M3>>", "M4": "<<woman matched with M4>>", "M5": "<<woman matched with M5>>" } <</answer>> Make sure that each man/woman is matched with exactly ONE partner. It is mandatory that you provide a matching as a JSON object enclosed in <<answer>><</answer>> tags as described above.
J.6 Example Prompt for Repeated Queries Due to Missing JSON Object
@@
For tasks where the desired output is a JSON object (when outputting a stable solution), models are given an additional opportunity to rectify issues in their response if the original response is incorrectly formatted. The prompt below is passed to the model to help rectify issues related to missing JSON objects. Note that the <<initially passed prompt>> and <<last 3,000 characters of LLM’s first response>> XML tags are replaced by the initial prompt and the tail of the models initial response, respectively.
@@
Previously, I gave you the following task: ——————————————————— <<initially passed prompt>> ——————————————————— In your response, you either failed to provide me with a matching or did not adhere to the JSON format I had asked for. Here are the last few lines of your response for reference: ——————————————————— <<last 3,000 characters of LLM’s first response>> ——————————————————— Please correct your response and provide me with the matching in the following JSON format, enclosed in <<answer>><</answer>> tags.<<answer>> { "M1": "<<woman matched with M1>>", "M2": "<<woman matched with M2>>", "M3": "<<woman matched with M3>>", "M4": "<<woman matched with M4>>", "M5": "<<woman matched with M5>>" } <</answer>> Make sure that each man/woman is matched with exactly ONE partner.
@@
J.7 Example Prompt for Repeated Queries Due to Incorrectly Formatted JSON Object
@@
Similar to the prompt in Section˜J.6, the following prompt is passed to the model when the model’s initial response contains an incorrectly formatted JSON object. Once again, the <<initially passed prompt>> and <<last 3,000 characters of LLM’s first response>> XML tags are replaced by the initial prompt and the tail of the models initial response, respectively.
@@
Previously, I gave you the following task: ——————————————————— <<initially passed prompt>> ——————————————————— In your response, you failed adhere to the JSON format I had asked for. Here are the last few lines of your response for reference: ——————————————————— <<last 3,000 characters of LLM’s first response>> ——————————————————— Please correct your response and provide me with the matching in the following JSON format, enclosed in <<answer>><</answer>> tags.<<answer>> { "M1": "<<woman matched with M1>>", "M2": "<<woman matched with M2>>", "M3": "<<woman matched with M3>>", "M4": "<<woman matched with M4>>", "M5": "<<woman matched with M5>>" } <</answer>> Make sure that each man/woman is matched with exactly ONE partner.
J.8 Example Prompt for Repeated Queries Due to Incomplete Matching
@@
Similar to the prompt in Section˜J.6, the following prompt is passed to the model when the model’s initial response contains a correctly formatted JSON object, but the matching itself is incomplete, or some agents have multiple partners. After the initially passed prompt, note that additional details are provided to assist the LLM in rectifying its response. Again, the <<initially passed prompt>> and <<last 3,000 characters of LLM’s first response>> XML tags are replaced by the initial prompt and the tail of the models initial response, respectively.
@@
Previously, I gave you the following task: ——————————————————— <<initially passed prompt>> ——————————————————— In your response, the matching you selected involves some women being matched with multiple men, which is not allowed. For example, W2 is matched with M1, M2, and M5. Additionally, W3, and W4 are unmatched. Here are the last few lines of your response for reference: ——————————————————— <<last 3,000 characters of LLM’s first response>> ——————————————————— Please correct your response and provide me with the matching in the following JSON format, enclosed in <<answer>><</answer>> tags.<<answer>> { "M1": "<<woman matched with M1>>", "M2": "<<woman matched with M2>>", "M3": "<<woman matched with M3>>", "M4": "<<woman matched with M4>>", "M5": "<<woman matched with M5>>" } <</answer>> Make sure that each man/woman is matched with exactly ONE partner.
Appendix K Training Examples for Fine-tuning
K.1 System-prompt ($\boldsymbol{s}$)
@@
K.2 High-level instruction ($\boldsymbol{u}$)
@@
K.3 Preference Profile ($\boldsymbol{p^{(i)}}$)
<<preferences>> { M: { M1: [W5,W3,W4,W2,W1], M2: [W2,W3,W5,W1,W4], M3: [W5,W3,W1,W4,W2], M4: [W1,W3,W2,W5,W4], M5: [W2,W3,W4,W1,W5], }, W: { W1: [M3,M5,M4,M1,M2], W2: [M1,M3,M4,M5,M2], W3: [M3,M2,M4,M1,M5], W4: [M4,M2,M3,M5,M1], W5: [M2,M4,M5,M1,M3], }} <</preferences>> <</preferences>>
K.4 Task-prompt ($\boldsymbol{t^{(i)}}$)
@@
    Your task is to find the proposer-optimal stable matching. Once you have found a stable matching, please return your matching in the JSON format given below: <<answer>> { "M1": "<<woman matched with M1>>", "M2": "<<woman matched with M2>>", "M3": "<<woman matched with M3>>", "M4": "<<woman matched with M4>>", "M5": "<<woman matched with M5>>" } <</answer>> Make sure that each man/woman is matched with exactly ONE partner. It is important that you enclose your JSON object in <<answer>><</answer>> tags.
@@
    Now, you will be asked a specific question about agent preferences: <<question>> Who is agent W3’s, 5-most preferred agent? <</question>> Once you have determined the answer, provide your output in the following format: 1. The solution as a single agent name. For example, "W1" Present your final answer within <<answer>> tags. IMPORTANT: ONLY RETURN THE NAME OF THE SINGLE AGENT THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
@@
    Now, you will be asked a specific question about agent preferences: <<question>> Would agent W3, prefer M4 and M3 over M1? <</question>> Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within <<answer>> tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
@@
    Now, you will be asked a specific question about agent preferences: <<question>> If agent W5 is currently engaged to M3, would she accept proposals from M1 and M4? <</question>> Once you have determined the answer, provide your output in the following format: 1. The solution as a YES or a NO. For example, "NO" Present your final answer within <<answer>> tags. IMPORTANT: ONLY RETURN YES OR NO THAT IS THE ANSWER TO THE QUESTION. Do not include any explanations or additional information in your final answer.
@@
### Response: <<think>>
K.5 Reasoning-trace ($\boldsymbol{r^{(i)}}$):
@@
    ### Response: <<think>> Okay, so I need to find a stable matching for 5 men and 5 women. For this, I can use the Gale-Shapley algorithm. Before I compute the solution for this instance, I will list the steps of the Gale-Shapley algorithm. 1. Initialize all men and women as unmatched. 2. Create a list to keep track of each man’s next proposal (initially set to 0 for all men). 3. While there are unmatched men: a. Select an unmatched man (M). b. Find the next woman (W) on M’s preference list that he hasn’t proposed to yet. c. If W is unmatched, match M and W. d. If W is matched but prefers M to her current partner: - Unmatch W from her current partner. - Match M and W. - Set the unmatched man as W’s previous partner. e. If W rejects M, move to the next woman on M’s preference list. 4. Repeat step 3 until all men are matched. Next, I will exectute the above algorithm for the given instance. The steps are as follows: M1 is free. M1 proposes to W3 Since W3 is free, W3 accepts the proposal. Now M1 and W3 are matched. M2 is free. M2 proposes to W5 Since W5 is free, W5 accepts the proposal. Now M2 and W5 are matched. M3 is free. M3 proposes to W1 Since W1 is free, W1 accepts the proposal. Now M3 and W1 are matched. M4 is free. M4 proposes to W4 Since W4 is free, W4 accepts the proposal. Now M4 and W4 are matched. M5 is free. M5 proposes to W4 Since W4 prefers their current partner M4 to M5, W4 rejects the proposal. M4 and W4 are still matched, and M5 is still free. M5 is free. M5 proposes to W1 Since W1 prefers M5 to their current partner M3, W1 accepts the proposal. Now M5 and W1 are matched, and M3 is free. M3 is free. M3 proposes to W4 Since W4 prefers their current partner M4 to M3, W4 rejects the proposal. M4 and W4 are still matched, and M3 is still free. M3 is free. M3 proposes to W3 Since W3 prefers their current partner M1 to M3, W3 rejects the proposal. M1 and W3 are still matched, and M3 is still free. M3 is free. M3 proposes to W2 Since W2 is free, W2 accepts the proposal. Now M3 and W2 are matched. <</think>>
@@
    Alright, I have been given the task of determining W3’s 5-most preferred agent. For this, let me look at her preference list in order: position 1: M4 position 2: M5 position 3: M1 position 4: M3 position 5: M2 Hence, W3’s 5-most preferred agent is M2. <</think>>